An engine absorbs 1.70kJ from a hot reservoir at 277oC and expels 1.20kJ to a cold reservoir at 27oC in each cycle.
(a) what is the engine's efficiency
(b) how much work is done by the engine in each cycle?
(c) what is the power output of the engine if each cycle lasts 0.300s?
1
Expert's answer
2021-03-01T12:41:23-0500

a)

"\eta=1-\frac{Q_2}{Q_1}=1-\frac{1200}{1700}\approx 0.294=29.4\%."

b)

"\Delta Q=Q_1-Q_2=1700-1200=500~J."

c)

"P=\frac{\Delta Q}{\Delta t}=\frac{500}{0.3}=1667~W."